# ERPoT: Effective and Reliable Pose Tracking for Mobile Robots Based on Lightweight and Compact Polygon Maps

### v1

@@ III-D2 Point-polygon matching
Given the polygon map ${\cal M}_{p}$, before commencing with pose tracking, the process of map initialization is carried out, and this step is crucial for establishing a solid foundation for accurate tracking. Specifically, for the $i$-th polygon $\mathbf{P}_{i}=\{\mathbf{p}_{1}^{w},\mathbf{p}_{2}^{w},...,\mathbf{p}_{N_{i}}^{w}\}$, the corresponding centroid point $\mathbf{p}_{i}^{e}=[x_{i}^{e},y_{i}^{e}]$ is calculated as
$x_{i}^{e}=\frac{m_{10}}{m_{00}},\ y_{i}^{e}=\frac{m_{01}}{m_{00}},$ (17)
Upon completion of the preliminary map initialization, the point-polygon matching process is executed, leveraging two point clouds denoted as ($\mathbb{C}_{e}$ and $\mathbb{C}_{v}$), as well as two sets of feature points ($\mathbb{F}_{c}$ and $\mathbb{F}_{e}$). Specifically, the polygon map form utilized in this study is actually a lightweight and accurate description of environmental occupancy information. Consequently, during the processes of feature matching or pose estimation, it is essential that the feature points within the current 2D scan align with or closely conform to the polygons as much as possible.
@@
Point-to-Vertex: In the realm of point-to-vertex matching, our methodology commences by constructing a KD-tree based on $\mathbb{C}_{v}$. Subsequently, assuming that the closest vertex retrieved from the KD-tree represents the accurate correspondence for each corner feature point.
@@
Point-to-Edge: For the purpose of point-to-edge matching, a KD-tree is meticulously constructed from the point cloud denoted as $\mathbb{C}_{e}$, where each point within $\mathbb{C}_{e}$ corresponds to a specific polygon. This KD-tree is subsequently employed to identify the $K$-nearest polygons associated with each edge feature point. In scenarios where a point is determined to be enclosed by the $K$-nearest polygons, as shown by ${\mathbf{q}}_{k_{3}}^{e}$ in Fig. 9, it is deliberately excluded from the subsequent phase of formulating the cost function.
@@ III-D3 Real-time pose tracking
Therefore, the optimal pose estimation can be obtained according to equation (18) based on current pose guess $\tilde{\mathbf{s}}_{t}$, current observation, and the prior polygon map.
@@ IV Experiments
To show the superiority and effectiveness of the proposed ERPoT, both self-recorded datasets and publicly available datasets are used for the evaluation in this section. As shown in Fig. 10, the experimental environment exemplifies a typical campus setting, characterized by an abundance of architectural structures and lush vegetation. On the other hand, publicly available datasets used in this evaluation include the KITTI dataset [46] and the Newer College dataset [47], and corresponding environments and platforms are shown in Fig. 11.
@@
Furthermore, Robot Operating System (ROS) is used for hardware and software implementations, and the codes used for polygon map construction and pose tracking are programmed in C/C++. The approach has been benchmarked against the other five approaches. During the pose tracking process, it is important to note that estimation of the current pose relies solely on data obtained from the LiDAR sensor.
@@
Figure 11: The experimental environments and the platforms for dataset collection of two publicly available datasets. For the KITTI dataset, the LiDAR sensor HDL-64E (Velodyne) is used to capture the corresponding point cloud dataset with typical street and road environments. The Newer College dataset is acquired using the handheld LiDAR sensor os1-64 (Ouster) at typical walking speeds through New College, Oxford for nearly $6.7$ km.
@@ IV-A Experimental configurations
(d) KITTI06
(e) KITTI08
Figure 12: The experimental configurations of the KITTI dataset. The blue thin line represents the complete trajectory of the dataset, and the black thick line represents the trajectory used to construct the prior map. In addition, the trajectory represented by the red thick line is used for experimental evaluation of pose tracking, and the corresponding start pose is marked by the magenta pentagram.
In the pursuit of pose tracking evaluation, we meticulously select segments from each dataset to construct prior maps, which act as essential baselines for the localization process. Within this subsection, we delve into a comprehensive description of the experimental configurations for each dataset, providing detailed insights into their respective setups.
Specifically, for the KITTI dataset, sequences 00, 02, 05, 06, and 08 have been deliberately selected for our evaluative purposes, leveraging their overlapping trajectories to ensure a reliable and comprehensive assessment of pose tracking performance. On the other hand, it is crucial to recognize that although the KITTI dataset remains the famous vision benchmark, the inconsistencies in the baselines are on the order of meters since the ground truth system relies only on RTK-GPS measurements [26]. Therefore, we combine point cloud registration with the ground truth provided by the KITTI dataset to further optimize the ground truth of the dataset for subsequent performance evaluation. The detailed experimental configurations are shown in Fig. 12, and the corresponding LiDAR sequences are shown in Tab. I.
TABLE I: Experimental configurations of KITTI dataset
@@ IV-B Qualitative experimental results
In this section, the qualitative experimental results of pose tracking based on the prior polygon map are provided. At first, based on the experimental configurations for self-recorded datasets and publicly available datasets, the corresponding prior polygon maps for the proposed ERPoT are obtained. Subsequently, the pose tracking process based on these prior maps is executed and illustrated in Fig. 15.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
For Self dataset-01, the process of pose tracking is represented by the three left subfigures in Fig. 15(a), and the obtained trajectory is smooth. As shown in the right subfigure in Fig. 15(a), the obtained 2D scan information is enough to describe the surrounding environmental information, including vegetation and buildings, which has a good matching relationship with the prior polygon map. Furthermore, within the Self dataset-02 environment, which includes uphill and downhill terrains, the proposed approach, while estimating only the 3-DoF $(x,y,\theta)$ of the pose, can still accurately estimate the pose during uphill and downhill movements, as demonstrated by the smooth trajectory in the Fig. 15(b).
On the other hand, for the KITTI00 dataset, our proposed ERPoT can realize effective and reliable pose tracking, even in scenarios where the collection platform navigates at high velocities within road environments. This capability is evident from the smooth trajectory illustrated in Fig. 15(c). In addition, in Fig. 15(d), for the handheld data collection platform, the smooth pose estimation result can also be obtained.
Drawing from the aforementioned qualitative experimental outcomes, the proposed approach ERPoT has demonstrated its capacity to achieve effective and reliable pose tracking across various data collection platforms equipped with a range of LiDAR sensors in different environments.
@@ IV-C Experimental evaluation metrics
Before we proceed with a detailed comparative quantitative evaluation, an overview of the qualitative measures employed for assessing performance is presented in this section. To evaluate the performance quality of the proposed effective and reliable pose tracking process, which is based on the lightweight and compact polygon map, a comprehensive set of performance metrics is employed.
Translation error and rotation error: For pose tracking that leverages a prior map, the system is endowed with a global reference, comprising a global position and a yaw angle. Consequently, the accuracy of pose estimation can be appraised using the Absolute Pose Error (APE) [48]. In this study, the $i$-th estimated pose is denoted by $\mathbf{{s}}_{i}=\left[{{{x}_{i}},{{y}_{i}},{{\theta}_{i}}}\right]$, and the ground truth is $\mathbf{\hat{s}}_{i}=\left[{{{\hat{x}}_{i}},{{\hat{y}}_{i}},{{\hat{\theta}}_{i}}}\right]$, then the error between $\mathbf{\hat{s}}_{i}$ and $\mathbf{s}_{i}$ can be parameterized as $\Delta\mathbf{s}_{i}=\left[{\Delta{x_{i}},\Delta{y_{i}},\Delta{\theta_{i}}}\right]$, and satisfies the following equation
@@
Runtime of pose tracking: In the realm of approach performance evaluation, the runtime serves as a pivotal metric, reflecting the efficiency of algorithms in processing information in various environments. The assessment of runtime is crucial for understanding the practical applicability of an approach, especially in time-sensitive environments. In this study, we incorporate the analysis of runtime as a key performance indicator, providing insights into the speed and responsiveness of our proposed ERPoT.
Size of the prior map: More importantly, the prior map, serving as a foundational element for accurate pose tracking, is meticulously designed to balance the trade-off between detail and efficiency, especially in large-scale environments with resource-constrained conditions. We advocate for a map that is not only rich in essential features but also optimized for minimal storage and swift retrieval, thereby facilitating real-time pose estimation without compromising computational resources. Accordingly, in this study, we emphasize the significance of a lightweight and compact prior map for pose tracking, which is a critical evaluation metric.
@@ IV-D Quantitative comparative results using public datasets
Based on the above experimental configurations and metrics, the comparative experiments between the proposed ERPoT and the other five approaches are carried out. First of all, we make use of the LiDAR sequences shown in Tab. I to construct the prior polygon map and the prior point cloud map, applied in the process of pose tracking. In addition, for SM_MCL, to efficiently perform the pose tracking, the distance field representation is used, which should be generated with a huge size in advance based on the point cloud map. Instead, the polygon maps for KITTI 00, 02, 05, 06, and 08 are shown in Fig. 16, which are compact and lightweight enough, ensuring that they encapsulate the most relevant spatial cues while excluding superfluous details that can impede processing speed. This design of polygon map representation enables our pose tracking approach to operate within a lean memory footprint, which is particularly advantageous for applications where hardware constraints are pronounced.
TABLE II: Experimental results of different approaches using the KITTI dataset
@@
SM_MCL $6709.49$ $99.72$ $33.00$ $38.15$ $2.07$ $0.55$ $0.66$ $68.03$ $\mathbf{36.09}^{\star}$ $\mathbf{5.60}^{\star}$
ERPoT(ours) $\mathbf{0.28}^{\star}$ $\mathbf{47.96}^{\star}$ $\mathbf{13.14}^{\star}$ $\mathbf{16.33}^{\star}$ $\mathbf{0.79}^{\star}$ $\mathbf{0.38}^{\star}$ $\mathbf{0.43}^{\star}$ $100.65$ $47.44$ $8.61$
@@
ERPoT(ours) $\mathbf{0.28}^{\star}$ $\mathbf{23.66}^{\star}$ $8.36$ $\mathbf{9.53}^{\star}$ $\mathbf{0.78}^{\star}$ $\mathbf{0.21}^{\star}$ $\mathbf{0.26}^{\star}$ $61.55$ $40.81$ $5.67$
@@
SM_MCL $6709.49$ $51.84$ $19.46$ $22.20$ $0.79$ $0.29$ $0.33$ $\mathbf{59.71}^{\star}$ $\mathbf{38.39}^{\star}$ $\mathbf{5.68}^{\star}$
ERPoT(ours) $\mathbf{0.28}^{\star}$ $\mathbf{34.46}^{\star}$ $12.88$ $16.64$ $\mathbf{0.41}^{\star}$ $\mathbf{0.12}^{\star}$ $\mathbf{0.15}^{\star}$ $61.71$ $45.28$ $6.64$
@@
ERPoT(ours) $\mathbf{0.85}^{\star}$ $\mathbf{33.13}^{\star}$ $\mathbf{11.58}^{\star}$ $\mathbf{13.39}^{\star}$ $\mathbf{0.64}^{\star}$ $\mathbf{0.23}^{\star}$ $\mathbf{0.26}^{\star}$ $68.50$ $43.83$ $\mathbf{6.84}^{\star}$
@@
ERPoT(ours) $\mathbf{0.85}^{\star}$ $\mathbf{83.44}^{\star}$ $\mathbf{18.70}^{\star}$ $\mathbf{22.10}^{\star}$ $\mathbf{0.85}^{\star}$ $\mathbf{0.29}^{\star}$ $\mathbf{0.35}^{\star}$ $\mathbf{74.72}^{\star}$ $48.40$ $\mathbf{8.01}^{\star}$
@@
SM_MCL $6647.48$ $44.22$ $15.55$ $17.78$ $\mathbf{1.52}$ $0.35$ $0.41$ $\mathbf{55.58}^{\star}$ $\mathbf{34.17}^{\star}$ $5.05$
ERPoT(ours) $\mathbf{0.36}^{\star}$ $\mathbf{25.46}^{\star}$ $\mathbf{7.94}^{\star}$ $\mathbf{9.01}^{\star}$ $1.58$ $\mathbf{0.24}^{\star}$ $\mathbf{0.33}^{\star}$ $58.91$ $42.63$ $\mathbf{4.71}^{\star}$
@@
ERPoT(ours) $\mathbf{0.36}^{\star}$ $\mathbf{26.84}^{\star}$ $\mathbf{10.14}^{\star}$ $\mathbf{11.14}^{\star}$ $\mathbf{0.60}^{\star}$ $\mathbf{0.26}^{\star}$ $\mathbf{0.29}^{\star}$ $\mathbf{68.72}^{\star}$ $\mathbf{41.59}^{\star}$ $\mathbf{5.97}^{\star}$
@@
ERPoT(ours) $\mathbf{0.27}^{\star}$ $\mathbf{28.78}^{\star}$ $\mathbf{7.25}^{\star}$ $\mathbf{8.08}^{\star}$ $\mathbf{0.51}^{\star}$ $\mathbf{0.23}^{\star}$ $\mathbf{0.24}^{\star}$ $\mathbf{63.79}^{\star}$ $46.55$ $\mathbf{5.63}^{\star}$
@@
ERPoT(ours) $\mathbf{0.59}^{\star}$ $78.01$ $\mathbf{16.10}^{\star}$ $\mathbf{22.47}^{\star}$ $\mathbf{0.96}^{\star}$ $\mathbf{0.34}^{\star}$ $\mathbf{0.40}^{\star}$ $67.14$ $45.39$ $5.33$
Furthermore, based on the obtained prior map, the process of pose tracking is conducted, and the pose tracking effect of ERPoT for KITTI00 is shown in Fig. 15(c). More importantly, Tab. II shows the quantitative comparative results based on the KITTI dataset, and the optimal result has been represented in bold font, while the experimental results for failed cases are indicated by ‘--’. It should be noted that for each test sequence, the success of pose tracking is confirmed when the final pose error remains within the specified limits—both the positional error within $200\mathrm{cm}$ and the angular error within $5^{\circ}$. This criterion ensures that only the sequences that meet these standards are subjected to the further evaluation.
As described in Tab. II, the proposed ERPoT successfully achieved pose tracking on eight test sequences. In particular, the innovative approach we propose makes use of a novel polygon-based prior map representation, which ensures compact and lightweight map storage. Across all five test environments, the map size is consistently under $1$ MB, demonstrating the efficiency and practicality of our method. Furthermore, regarding translational and rotational errors, the proposed approach has also demonstrated optimal performance. The mean values for translational and rotational errors are both below $20\mathrm{cm}$ and $0.5^{\circ}$, respectively, across all test scenarios, underscoring the accuracy and precision of our method. In addition, ERPoT demonstrates exceptional runtime performance, reliably achieving a consistent processing time of approximately $40\mathrm{ms}$ with high stability.
Refer to captionRefer to caption
@@
TABLE III: Experimental results of different approaches using the Newer College dataset
@@
Newer College-1 ALOAM_MCL 62.4962.49{62.49}62.49 $122.38$ $36.10$ $42.90$ $11.31$ $1.98$ $2.48$ $639.64$ $182.60$ $143.01$

### v3

@@ III-D2 Point-polygon matching
Based on the prior map obtained through polygon map construction, the point-polygon matching process is executed in this subsection, leveraging two point clouds denoted as ($\mathbb{C}_{e}$ and $\mathbb{C}_{v}$), as well as two sets of feature points ($\mathbb{F}_{c}$ and $\mathbb{F}_{e}$). In addition, the polygon map ${\cal M}_{p}$ with counter-clockwise polygon vertices is also used for the point-to-edge matching step. Note that before implementing a matching process, it is necessary to ensure that the current number of observed corner features and edge features are greater than the thresholds $\tau_{c}$ and $\tau_{e}$, respectively.
Specifically, the polygon map form utilized in this study is actually a lightweight and accurate description of environmental occupancy information. Consequently, during the processes of feature matching or pose estimation, it is essential that the feature points within the current 2D scan align with or closely conform to the polygons as much as possible.
@@
Point-to-Vertex: In the realm of point-to-vertex matching, our methodology commences by constructing a KD-tree based on $\mathbb{C}_{v}$. Subsequently, assuming that the closest vertex retrieved from the KD-tree represents the accurate correspondence for each corner feature point, provided that the distance from the closest vertex is less than the given threshold $\epsilon_{v}$.
@@
Point-to-Edge: For the purpose of point-to-edge matching, a KD-tree is meticulously constructed from the point cloud denoted as $\mathbb{C}_{e}$, where each point within $\mathbb{C}_{e}$ corresponds to a specific polygon in ${\cal M}_{p}$ with counter-clockwise vertices. This KD-tree is subsequently employed to identify the $K$-nearest polygons associated with each edge feature point. In addition, the distance from the closest polygon should be less than the given threshold $\epsilon_{p}$. In scenarios where a point is determined to be enclosed by the $K$-nearest polygons, as shown by ${\mathbf{q}}_{k_{3}}^{e}$ in Fig. 9, it is deliberately excluded from the subsequent phase of formulating the cost function.
@@ III-D3 Real-time pose tracking
Therefore, the optimal pose estimation can be obtained according to equation (18) based on current pose guess $\tilde{\mathbf{s}}_{t}$, current observation, and the prior polygon map. Subsequently, the reliability of the current tracking result is assessed based on the discrepancy between the solved pose information and the predicted pose information. At the same time, the distance and angle thresholds are set to $\lambda_{d}$ and $\lambda_{a}$, respectively.
@@ IV Experiments
To show the superiority and effectiveness of the proposed ERPoT, both self-recorded datasets and publicly available datasets are used for the evaluation in this section. As shown in Fig. 10, the experimental environment exemplifies a typical campus setting, characterized by an abundance of architectural structures and lush vegetation. Moreover, publicly available datasets used in this evaluation include the KITTI dataset [46], the Newer College dataset [47], and the FusionPortable dataset [48], while corresponding environments and platforms are shown in Fig. 11.
@@
TABLE I: Hardware and software specification
Robot 204AKZ and FR-09
Sensor VLP-32C and Pandar40P
Processor Intel Core i7-11800H @2.3Ghz
RAM 32GB
Operating system Ubuntu 20.04
ROS Noetic
TABLE II: Systems parameters
Parameter: value
Grid map resolution ($r$)
- Small scenes $0.05$m
- Large scenes $0.10$m
Pose tracking
- Vertex data association threshold ($\epsilon_{v}$) $2.0$m
- Polygon data association threshold ($\epsilon_{p}$) $3.0$m
- Distance difference threshold ($\lambda_{d}$) $1.0$m
- Angle difference threshold ($\lambda_{a}$) $0.2$rad
Feature extraction
- Corner smoothness threshold ($\sigma_{c}$) $10.0$
- Edge smoothness threshold ($\sigma_{e}$) $0.1$
- Corner feature number threshold ($\tau_{c}$) $10$
- Edge feature number threshold ($\tau_{e}$) $50$
Furthermore, the Robot Operating System (ROS) is utilized for both hardware and software implementations. The code for constructing the polygon map and performing pose tracking is written in C/C++. Experiments are conducted on a computer with specifications detailed in Tab. I and the system parameters of ERPoT are shown in Tab. II. The approach has been benchmarked against six other methods. It is important to note that during the pose tracking process, the estimation of the current pose relies solely on the prior map (including the point cloud map, mesh map, distance field map, and the proposed polygon map) and data obtained from the LiDAR sensor.
@@
(c) FusionPortable dataset
Figure 11: The experimental environments and platforms for dataset collection of two publicly available datasets. For the KITTI dataset, the LiDAR sensor HDL-64E (Velodyne) is used to capture the corresponding point cloud dataset with typical street and road environments. The Newer College dataset is acquired using the handheld LiDAR sensor os1-64 (Ouster) at typical walking speeds through New College, Oxford for nearly $6.7$ km. Additionally, the FusionPortable dataset, captured using Ouster os1-128, is a typical campus-scene dataset that includes both indoor and outdoor environments.
@@
Range_MCL: The range image provides a lightweight 3D LiDAR scan representation, and the mesh map is more compact than the point cloud map. In [19], a novel sensor model combined with MCL is used for 3D LiDAR-based global localization and pose tracking  [^8] https://github.com/PRBonn/range-mcl. This model compares current LiDAR range images to synthetic ones from the mesh to adjust particle weights, offering a simple yet versatile approach.
@@ IV-A Experimental configurations
In the pursuit of pose tracking evaluation, we meticulously select segments from each dataset to construct prior maps, which act as essential baselines for the localization process. Specifically, we make use of the provided segment sequence including ground truth values and corresponding point clouds to transform all point clouds into the world coordinate system, thereby creating a prior point cloud map. Additionally, to manage the large scale of the point cloud map, we apply voxel filtering to facilitate efficient online pose tracking, the final prior map is used for ALOAM_MCL, FLOAM_MCL, KISS_MCL, and HDL_localization. Based on the obtained point cloud map, the distance field representation for SM_MCL is generated using the provided open-source tool. Similarly, the prior mesh map for Range_MCL is acquired using the open-source code.
@@
(d) KITTI08
(e) KITTI06
Figure 12: The experimental configurations of the KITTI dataset. The blue thin line represents the complete trajectory of the dataset, and the black thick line represents the trajectory used to construct the prior map. In addition, the trajectory denoted by the red thick line is used for evaluation of pose tracking, and the corresponding start pose is marked by the magenta pentagram.
For the KITTI dataset, sequences 00, 02, 05, 06, and 08 have been deliberately selected for our evaluative purposes, leveraging their overlapping trajectories to ensure a reliable and comprehensive assessment of pose tracking performance. On the other hand, it is crucial to recognize that although the KITTI dataset remains the famous vision benchmark, the inconsistencies in the baselines are on the order of meters since the ground truth system relies only on RTK-GPS measurements [26]. Therefore, we combine point cloud registration with the ground truth provided by the KITTI dataset to further optimize the ground truth of the dataset for subsequent performance evaluation. The detailed experimental configurations are shown in Fig. 12, and the corresponding LiDAR sequences are shown in Tab. III.
TABLE III: Experimental configurations of KITTI dataset
@@ 08 $0$-1306 1400-1850
(a) Portable-campus
(b) Portable-building
(c) Portable-garden
(d) Portable-canteen
Figure 14: The experimental configurations of the FusionPortable dataset.
For the FusionPortable dataset, the experimental configurations are illustrated in Fig. 14. Specifically, for the Portable-canteen and Portable-garden datasets, which encompass both day and night scenes, the night datasets (20220215_canteen_night and 20220215_garden_night) are used to construct the prior maps, while the day datasets (20220216_canteen_day and 20220216_garden_day) are employed for pose tracking evaluation. For Portable-campus (20220226_campus_road_day), the sequence of LiDAR frames within the range $[1,1100]$ is used for prior map construction, and the sequence $[11063,12263]$ is utilized for experimental evaluation. For Portable-building, which includes both indoor and outdoor scenes, the range $[1201,4500]$ is used for prior map construction, and the sequences $[1,1200]$ and $[4501,5996]$ are used for evaluation.
@@ IV-B Qualitative experimental results
In this section, the qualitative experimental results of pose tracking based on the prior polygon map are provided. At first, based on the experimental configurations for self-recorded datasets and publicly available datasets, the corresponding prior polygon maps for the proposed ERPoT are obtained. Subsequently, the pose tracking process based on these prior maps is carried out and illustrated in Fig. 16.
@@
For Self dataset-01, the process of pose tracking is represented by the three left subfigures in Fig. 16(a), and the obtained trajectory is smooth. As shown in the right subfigure in Fig. 16(a), the obtained 2D scan information is sufficient to describe the surrounding environmental information, including vegetation and buildings, which has a good matching relationship with the prior polygon map. Furthermore, within the Self dataset-02 environment, which includes uphill and downhill terrains, the proposed approach, while estimating only the 3-DoF $(x,y,\theta)$ of the pose, can still accurately estimate the pose during uphill and downhill movements, as demonstrated by the smooth trajectory in the Fig. 16(b).
On the other hand, for the KITTI00 dataset, our proposed ERPoT can realize effective and reliable pose tracking, even in scenarios where the collection platform navigates at high velocities within road environments. This capability is evident from the smooth trajectory illustrated in Fig. 16(c). In addition, in both Fig. 16(d) and Fig. 16(e), for the handheld data collection platform with different LiDAR sensors, the smooth pose estimation result can also be obtained.
Drawing from the aforementioned qualitative experimental outcomes, the proposed approach ERPoT has demonstrated its capacity to achieve effective and reliable pose tracking.
@@ IV-C Experimental evaluation metrics
Before the detailed evaluation, this section provides an overview of the qualitative measures used to assess performance. A comprehensive set of metrics evaluates the pose tracking process based on a lightweight polygon map.
Translation error and rotation error: For pose tracking using a prior map, the system has a global reference with a global position and yaw angle. The accuracy of pose estimation is evaluated using the Absolute Pose Error (APE) [49]. In this study, the $i$-th estimated pose is denoted by $\mathbf{{s}}_{i}=\left[{{{x}_{i}},{{y}_{i}},{{\theta}_{i}}}\right]$, and the ground truth is $\mathbf{\hat{s}}_{i}=\left[{{{\hat{x}}_{i}},{{\hat{y}}_{i}},{{\hat{\theta}}_{i}}}\right]$, then the error between $\mathbf{\hat{s}}_{i}$ and $\mathbf{s}_{i}$ can be parameterized as $\Delta\mathbf{s}_{i}=\left[{\Delta{x_{i}},\Delta{y_{i}},\Delta{\theta_{i}}}\right]$, and satisfies the following equation
@@
Runtime of pose tracking: In the realm of approach performance evaluation, runtime is a crucial metric, reflecting the efficiency of algorithms in processing information in various environments. Assessing runtime is essential for understanding the practical applicability of an approach, especially in time-sensitive environments. In comparative experiments, we report only CPU times, focusing on CPU performance to ensure consistency across different algorithms. The times cover the full pose tracking cycle, including data preprocessing, feature extraction, feature matching, and pose estimation.
Size of the prior map: Moreover, as a foundational element for accurate pose tracking, the prior map is designed to balance detail and efficiency, especially in large-scale, resource-constrained environments. We advocate for a map that is rich in essential features yet optimized for minimal storage and swift retrieval, facilitating real-time pose estimation without compromising computational resources.
@@ IV-D Quantitative comparative results using public datasets
Based on the experimental configurations and metrics, comparative experiments are conducted between ERPoT and six other approaches. The prior polygon and point cloud maps are constructed using the LiDAR sequences listed in Tab. III for pose tracking. For SM_MCL, a distance field representation, which requires significant precomputation based on the point cloud map, is used for efficient pose tracking. In contrast, the polygon maps for KITTI sequences 00, 02, 05, 06, and 08, shown in Figure 17, are lightweight and compact. They capture essential spatial cues while omitting unnecessary details. This design allows ERPoT to operate with minimal memory usage, which is crucial for applications with hardware limitations.
TABLE IV: Experimental results of different approaches using the KITTI dataset
@@
SM_MCL $6709.49$ $99.72$ $33.00$ $38.15$ $2.07$ $0.55$ $0.66$ $68.03$ $\mathbf{36.09}^{\star}$ 5.605.60{5.60}5.60
Range_MCL 12.6712.67{12.67}12.67 100.11100.11{100.11}100.11 73.3173.31{73.31}73.31 74.7674.76{74.76}74.76 6.076.07{6.07}6.07 2.402.40{2.40}2.40 2.832.83{2.83}2.83 $135.73$ $132.13$ $\mathbf{4.59}^{\star}$
ERPoT(ours) $\mathbf{0.28}^{\star}(2584/34708)$ $\mathbf{47.96}^{\star}$ $\mathbf{13.14}^{\star}$ $\mathbf{16.33}^{\star}$ $\mathbf{0.79}^{\star}$ $\mathbf{0.38}^{\star}$ $\mathbf{0.43}^{\star}$ $100.65$ $47.44$ $8.61$
@@
Range_MCL ${12.67}$- -- -- -- -- -- -- -- --
ERPoT(ours) $\mathbf{0.28}^{\star}(2584/34708)$ $\mathbf{23.66}^{\star}$ $8.36$ $\mathbf{9.53}^{\star}$ $\mathbf{0.78}^{\star}$ $\mathbf{0.21}^{\star}$ $\mathbf{0.26}^{\star}$ $61.55$ $40.81$ $5.67$
@@
SM_MCL $6709.49$ $51.84$ $19.46$ $22.20$ $0.79$ $0.29$ $0.33$ $\mathbf{59.71}^{\star}$ $\mathbf{38.39}^{\star}$ $5.68$
Range_MCL 12.6712.67{12.67}12.67 152.38152.38{152.38}152.38 112.33112.33{112.33}112.33 118.33118.33{118.33}118.33 3.663.66{3.66}3.66 1.091.09{1.09}1.09 1.391.39{1.39}1.39 $135.92$ $132.38$ $\mathbf{1.72}^{\star}$
ERPoT(ours) $\mathbf{0.28}^{\star}(2584/34708)$ $\mathbf{34.46}^{\star}$ $12.88$ $16.64$ $\mathbf{0.41}^{\star}$ $\mathbf{0.12}^{\star}$ $\mathbf{0.15}^{\star}$ $61.71$ $45.28$ $6.64$
@@
Range_MCL ${27.38}$- -- -- -- -- -- -- -- --
ERPoT(ours) $\mathbf{0.85}^{\star}(7582/106065)$ $\mathbf{33.13}^{\star}$ $\mathbf{11.58}^{\star}$ $\mathbf{13.39}^{\star}$ $\mathbf{0.64}^{\star}$ $\mathbf{0.23}^{\star}$ $\mathbf{0.26}^{\star}$ $68.50$ $43.83$ $\mathbf{6.84}^{\star}$
@@
Range_MCL ${27.38}$- -- -- -- -- -- -- -- --
ERPoT(ours) $\mathbf{0.85}^{\star}(7582/106065)$ $\mathbf{83.44}^{\star}$ $\mathbf{18.70}^{\star}$ $\mathbf{22.10}^{\star}$ $\mathbf{0.85}^{\star}$ $\mathbf{0.29}^{\star}$ $\mathbf{0.35}^{\star}$ $\mathbf{74.72}^{\star}$ $48.40$ $\mathbf{8.01}^{\star}$
@@
SM_MCL $6647.48$ $44.22$ $15.55$ $17.78$ $\mathbf{1.52}^{\star}$ $0.35$ $0.41$ $\mathbf{55.58}^{\star}$ $\mathbf{34.17}^{\star}$ $5.05$
Range_MCL ${23.37}$- -- -- -- -- -- -- -- --
ERPoT(ours) $\mathbf{0.36}^{\star}(3321/44174)$ $\mathbf{25.46}^{\star}$ $\mathbf{7.94}^{\star}$ $\mathbf{9.01}^{\star}$ $1.58$ $\mathbf{0.24}^{\star}$ $\mathbf{0.33}^{\star}$ $58.91$ $42.63$ $\mathbf{4.71}^{\star}$
@@
Range_MCL 23.3723.37{23.37}23.37 176.95176.95{176.95}176.95 83.4983.49{83.49}83.49 91.0591.05{91.05}91.05 5.805.80{5.80}5.80 1.321.32{1.32}1.32 1.661.66{1.66}1.66 $235.29$ $151.41$ $88.27$
ERPoT(ours) $\mathbf{0.36}^{\star}(3321/44174)$ $\mathbf{26.84}^{\star}$ $\mathbf{10.14}^{\star}$ $\mathbf{11.14}^{\star}$ $\mathbf{0.60}^{\star}$ $\mathbf{0.26}^{\star}$ $\mathbf{0.29}^{\star}$ $\mathbf{68.72}^{\star}$ $\mathbf{41.59}^{\star}$ $\mathbf{5.97}^{\star}$
@@
Range_MCL 10.8010.80{10.80}10.80 153.86153.86{153.86}153.86 110.38110.38{110.38}110.38 112.82112.82{112.82}112.82 2.892.89{2.89}2.89 0.750.75{0.75}0.75 0.950.95{0.95}0.95 $221.04$ $175.73$ $23.96$
ERPoT(ours) $\mathbf{0.27}^{\star}(2644/33504)$ $\mathbf{28.78}^{\star}$ $\mathbf{7.25}^{\star}$ $\mathbf{8.08}^{\star}$ $\mathbf{0.51}^{\star}$ $\mathbf{0.23}^{\star}$ $\mathbf{0.24}^{\star}$ $\mathbf{63.79}^{\star}$ $46.55$ $\mathbf{5.63}^{\star}$
@@
Range_MCL ${25.14}$- -- -- -- -- -- -- -- --
ERPoT(ours) $\mathbf{0.59}^{\star}(5451/71972)$ $78.01$ $\mathbf{16.10}^{\star}$ $\mathbf{22.47}^{\star}$ $\mathbf{0.96}^{\star}$ $\mathbf{0.34}^{\star}$ $\mathbf{0.40}^{\star}$ $67.14$ $45.39$ $5.33$
Furthermore, pose tracking is conducted using the prior map, and the effect for KITTI00 is shown in Fig. 16(c). More importantly, Tab. IV shows the quantitative comparative results on KITTI dataset, with optimal results in bold and failed cases indicated by ‘--’. During the pose tracking experiments for each test sequence, success is confirmed when the final pose error remains within the specified limits: translational error within $200\mathrm{cm}$ and rotational error within $20^{\circ}$. Additionally, the pose error during the process must be less than $500\mathrm{cm}$ and $30^{\circ}$. This criterion ensures that only the sequences meeting these standards are subjected to further evaluation.
As shown in Tab. IV, ERPoT successfully achieved pose tracking on nine test sequence, which uses a novel polygon-based prior map for compact and lightweight storage. Across all five test environments, the map size is consistently under $1$ MB, demonstrating the efficiency and practicality of our method, and the number of polygons/vertices is displayed after the map size. Furthermore, regarding translational and rotational errors, the proposed approach has also demonstrated optimal performance. The mean values for translational and rotational errors are both below $20\mathrm{cm}$ and $0.5^{\circ}$, respectively, across all test scenarios, underscoring the accuracy and precision of our method. In addition, ERPoT demonstrates exceptional runtime performance, reliably achieving a consistent processing time of approximately $40\mathrm{ms}$ with high stability.
@@
TABLE V: Experimental results of different approaches using the Newer College dataset
@@
Newer College-1 ALOAM_MCL 62.4962.49{62.49}62.49 $\mathbf{122.38}^{\star}$ $36.10$ $42.90$ $11.31$ $1.98$ $2.48$ $639.64$ $182.60$ $143.01$